# HOLMBERG IX: A UNIQUE, INFANT BUT INACTIVE GALAXY as Revealed via a Multiwavelength Approach

### 03 Apr 2026 v1

@@ III.2 Uniformly Young Stellar Population Demonstrated by UV Diagnosis
Stellar population age for Holmberg IX is diagnosed by analyzing the 4 UV images, GALEX FUV, NUV, and Swift-UVOT UVW2, UVW1, on a spatially resolved basis with the aid of stellar population synthesis modeling. The UV bands have special advantage in prospecting for young stellar populations, in contrast to previous optical and NIR studies (Sabbi et al., 2008; Weisz et al., 2008; Okamoto et al., 2015). The following sections present the UV diagnosis in both pictorial and graphic manners.
@@ III.2.2 Color-color Diagram
From the color-color diagram, we can see that the FUV clusters and the diffuse regions inside the elliptical aperture possess the same regime following the model curve well;999There are two clusters with behaving as the exception. They are separately located on the galactic periphery and suspected to be background sources. the main body of the galaxy as a whole and the H-emission area are positioned identically at mag and mag, corresponding to the age Myr by reference to the model, as an average of the subregions. The overlaps between the different types of regions in this diagram imply all of them indistinguishable in stellar population. Therefore, the FUV clusters are not definitely younger than the diffuse regions but merely more in the number of stars, so is the H-emission area. In this situation, all of stars in Holmberg IX are supposed to be the same population born from a single SF event Myr ago, as also depicted by the single-peak Gaussian profiles in Figure 2. The recent occurrence of the global star formation coincides with the hypothesis of the tidal origin for this galaxy. Combining with the above 2D map, we can demarcate for the galactic mainbody the lower age limit Myr that is newly discovered through the UV diagnosis in this work, and the upper limit Myr that is coincident with the previous optical-NIR studies ( Myr, Sabbi et al., 2008; Weisz et al., 2008; Okamoto et al., 2015). The lack of younger stellar populations suggests a very short duration for the past SF event. The model curve in Figure 3 is reproduced with the SFH of an instantaneous SF burst. Justification for the applied SFH is presented in Appendix A with more complex SFHs compared and discussed.
@@ III.3 Sparse Ionized Gas Revealed by H Mapping
For the purpose of exploring ionized gas linking current SF activities, we observed Holmberg IX through H channel in both of the direct narrowband imaging and FPI scanning modes (described in Section II.1). From a general view provided by the direct narrowband image, the H emission emerges dim, diffuse, and dispersive all over the galaxy. In the deep-exposure FPI map, it is spatially resolved into pieces of scattered spots and filaments located at north-west and south-east peripheries; most of the galactic body is left blank. The H appearance is opposite to the conspicuous and concentrative UV emission exhibited in the GALEX and Swift-UVOT images. Similar H-UV spatial contrast has been found in some other nearby dwarf galaxies and interpreted with SF propagation from galactic centers to outskirts, so that after a certain timescale, current SF with strong H arrives in rims surrounding recently formed stellar populations bright in UV (Egorov et al., 2018; Gerasimov et al., 2024; Yarovova et al., 2024). Notwithstanding, in this work such a behavior is observed in a TDG (candidate) for the first time.
@@
IV IMPLICATION
Under average conditions, weak H emission indicates a currently quiescent SF state. For Holmberg IX, this feature in combination with the UV diagnosis in Section III.2 potentially manifests global SF already ceased in most area of the galaxy Myr ago. In this section, we assess current star formation activities reflected by the marginal H plus the luxuriant FUV emissions, and deduce possible evolutionary pathways for this galaxy.
@@ IV.1 Current Star Formation State
The differences between the H and UV emissions in spatial position and radiative intensity inevitably cause considerable discrepancy between SFRs estimated from H and FUV integrated luminosities. By taking the aperture photometry and adopting the calibrations in Kennicutt (1998, hereafter denoted as K98), we obtain SFR(H) = () (corresponding to () ) and SFR(FUV) = () (corresponding to () ). Via specifically sampling the H-emission area, the selective photometry plus the K98 calibrations yields SFR(H) = () (corresponding to () ), SFR(FUV) = () (corresponding to () ), which narrows the gap between SFR(H) and SFR(FUV) effectively but remains insufficient. This examination demonstrates that the spatial mismatch between H and FUV emissions contributes a major part of the common discrepancy between SFR(H) and SFR(FUV) if galaxies are measured on a integrated basis. In this instance, different sources with H and FUV are actually sampled in practice. The selective photometry enables approaching the same H and FUV positions but still fails to totally fill the gap between the SFRs.
Different SF timescales traced by H and FUV luminosities are considered primarily responsible for SFR(H) lower than SFR(FUV), so-called ”H-to-FUV deficit”, no matter whether the two emissions are spatially displaced or identically located. From this perspective, SFR(H) takes an average of the latest Myr while SFR(FUV) averages past Myr (Kennicutt, 1998). In this case, SFR(FUV) higher than SFR(H) is a plausible consequence of different SF states tens of Myr ago (active) and at present (quiescent).
In order to quantify SF activity levels, we adopt the birthrate parameter defined as the ratio of current to mean past SFRs ( ; Kennicutt et al., 1994). The proportion valued by weighs the current star-formation level on the whole lifetime scale of a galaxy and reflects the evolution of SF activities. In this work, we set the mean past SFR equal to , where is total stellar mass and is the stellar population age.101010The standard definition of the birthrate parameter contains a fraction of stellar mass returned to the galaxy in a stellar generation (Kennicutt et al., 1994). In this work, we assume a zero return fraction of stellar mass. Therefore, all mass is newly generated in every SF event, so that the birthrate parameter reaches the maximum. and are obtained with the stellar population synthesis model as best-fit parameters. The numerator SFRcurrent in is assigned as SFR(H) and SFR(FUV), respectively, which consequently brings on the birthrate parameter representing the SF state in recent Myr by and that in recent Myr by .
@@ IV.2 Subsequent Evolution and Ultimate Fate
The global SF dearth albeit the overall infancy for stellar populations requires rapid SF quenching, for which the reason seems to be a conundrum, particularly in view that the galaxy resides in a gas-rich environment ( and for Holmberg IX, presented in Sabbi et al., 2008). At the same time, the currently quiescent state conflicts with commonly accepted requisite of continuous SF for TDGs (Duc & Mirabel, 1999, and references therein) and raises a question on subsequent evolution and ultimate fate of Holmberg IX. Without SF continuing, Holmberg IX is supposed to stay in a transient status at present and be transforming into a long-lived dwarf spheroidal galaxy due to gravitational contraction similar to those prevailing in the local group,111111In the local group, although there is no TDG, spheroidal galaxies are likely to originate from tidal events, which has been predicted by a number of numerical simulations (e.g., Mateo, 1998; Casas et al., 2012; Yang et al., 2014)., or to the opposite, dissolving into an ultra-diffuse galaxy (UDG) or intergalactic stellar stream because of centrifugal effects (and perhaps vanish in the universe eventually). At this present moment, we are probably witnessing a rarely captured phenomenon for a TDG.
@@ V DISCUSSION
The above assessment and deduction rely on a high-mass SF scenario which ascribes weak H but strong FUV emanations (or H-to-FUV deficit) to SFR low in recent Myr but high tens of Myr ago with bursty SF. In this case, aging of stellar populations leads to lack of high-mass, short-lifetime ionizing stars, which weakens H but sustains longer-time luminescence of FUV (Sullivan et al., 2000; Bell & Kennicutt, 2001; Sullivan et al., 2004; Iglesias-Páramo et al., 2004; Weisz et al., 2012). However, some observations of dwarf galaxies by resolving individual stars have suggested a more complicated situation that high-mass ionizing stars are not always absent for H-to-FUV deficit (Weisz et al., 2008; McQuinn et al., 2009, 2010a, 2010b). In practical studies, IMF is usually assumed in a universal form with massive SF dominating, but if it turns out to be variant in some circumstance, the speculation about galactic properties would be different. Besides IMF forms, a diversity of other factors including SF modalities, interstellar medium conditions, observational bias, etc. are also able to interpret H-to-FUV deficit recognized in a number of dwarf galaxies and extended outer disks (so-called extended UV or XUV disks) of spiral galaxies (e.g., Hassani et al., 2024, and references cited in this section).
Instead of currently quiescent SF activities, the alternative explanations of H-to-FUV deficit allow current occurrence of intense SF in agreement with the common expectation for TDGs. For unity of SFR(H) and SFR(FUV), Lee et al. (2009, hereafter denoted as L09) have recalibrated to SFR for dwarf galaxies with erg s-1. By applying the L09 calibration to Holmberg IX in our work, we obtain SFR(H) = () (corresponding to () ) for the aperture photometry, indeed reducing the offset from SFR(FUV), and SFR(H) = () (corresponding to () ) for the selective photometry, even exceeding SFR(FUV). Nevertheless, applicability of the L09 prescription to all dwarf galaxies is uncertain and the resultant current SF activity for Holmberg IX is questionable until solid evidence is acquired. In below subsections, we elucidate all the other attributions of H-to-FUV deficit and discuss the suitability for Holmberg IX.
@@ V.1 Low-mass or/and Stochastic SF
A low H amount is most directly ascribed to inherent shortage of high-mass ionizing stars. In addition to aging of stellar populations formed on a short or instantaneous timescale as suggested above, continuous SF with IMF in a low-mass dominant form is another mechanism innately preventing massive SF and consequently inducing H-to-FUV deficit. The variant IMF has been not only reproduced by stellar population synthesis modeling (Kroupa & Weidner, 2003; Weidner & Kroupa, 2005, 2006; Pflamm-Altenburg et al., 2007, 2009) but also observed in local galaxies with low surface brightness (Hoversten & Glazebrook, 2008; Meurer et al., 2009) and galactic XUV disks (Bruzzese et al., 2020; Rautio et al., 2024), still it is not regarded universally effective (Hunter et al., 2010).
Even if stars keep forming with invariantly normal IMF, lack of high-mass stars is also likely caused by stochasticity in SF events (da Silva et al., 2012, 2014). It is possible for random sampling of constant IMF at low-SFR to undersample the number of high-mass ionizing stars (Lee et al., 2009, and references therein), but the stochastic SF seems still not to suffice for the explanation either (Weisz et al., 2012). Therefore, two or more combinative effects of bursty, low-mass, or/and stochastic SF are considered more reasonable for restricting massive SF and more realistic for explaining observational H/FUV deficit (Fumagalli et al., 2011; da Silva et al., 2014; Lee et al., 2016). In this case, Holmberg IX is likely to keep forming stars for maintaining long-time survival, which hints there are possibly TDGs of the same type that failed to be recognized in the universe.
@@ V.2 Loss of Ionizing Photons
If high-mass ionizing stars are always sufficient, H/FUV deficit is likely caused by loss of ionizing photons. The most probable mechanism is leakage of Lyman continuum in thin, porous, or/and (semi)transparent interstellar medium which tends to decrease SFR(H) by up to 25% at fixed SFR(FUV) (Relaño et al., 2012). This mechanism has been applied to dwarf galaxies and XUV disks as an alternative explanation of H/FUV deficit (Hunter et al., 2010, 2011, 2013). However, it seems not possible for Holmberg IX located in a gas-rich environment where interstellar medium is rigorous and tight. Even though the HII regions are leaky and ionizing photons escape from natal areas, most of these photons still stay within the galaxy and supply diffuse emission lines which is accounted in integrated measurements.
@@ V.3 Underestimates of Existing H Emission
In addition to lack of ionizing stars and loss of ionizing photons, underestimates of H emission potentially contribute to the offset between observed and expected H levels, particularly for low surface-brightness systems (Hunter et al., 2010). However, much deeper imaging of dwarf galaxies than previous surveys has evinced that unaccounted part of H emission (if any) compensates for only 5% flux at maximum and thus fails to suffice for a self-contained explanation (Lee et al., 2016). Nay, our ultra-deep FPI observations (with the sensitivity limit lower than that in Lee et al., 2016) and the net emission-line extraction from the 3D datacube guarantee the completeness of sampling H emission and thereof prevent the underestimation.
@@ VI SUMMARY
Via new observations deeply mapping H emission in Holmberg IX and combinatory analysis of archival/published data, we reveal a series of unique characteristics for this TDG candidate, distinct from other TDGs ever known. In the results, the galaxy appears extremely dustless, disproportionate to its rich gas and metal, with globally young stellar populations but currently inactive SF reflected by H/FUV deficit. In the galactic mainbody, all stars are suggested to be born from a bursty SF event Myr ago which then rapidly ceased, demarcating the lower age limit Myr; younger stars reside in outskirts where mild SF remains as an aftermath of SF radial propagation. The recent quenching seems like a conundrum in view of the gas-rich location of the galaxy.
On normal massive-SF assumption, the present quiescence conflicts with common expectation for TDGs hosting continuous SF, which hints Holmberg IX currently staying in a rarely captured stage transforming into a dwarf spheroidal galaxy or oppositely into a(n) (ultra-)diffuse (intergalactic) system which will probably dissolve in the end. Alternatively, by assuming a peculiar form of IMF, it is also possible for Holmberg IX to keep forming low-mass stars with H/FUV deficit and hence be able to maintain the current status for long-time survival. In whichever situation in reality, differing from conventional understandings, Holmberg IX deserves further investigations of its past experience, present state, and future destiny, as well as explorations of analogous TDGs in the universe.

### 16 Apr 2026 v2

@@ III.2 Uniformly Young Stellar Population Demonstrated by UV Diagnosis
Stellar population age for Holmberg IX is diagnosed by analyzing the four UV images, GALEX FUV, NUV, and Swift-UVOT UVW2, UVW1, on a spatially resolved basis with the aid of stellar population synthesis modeling. The UV bands have special advantage in prospecting for young stellar populations, in contrast to previous optical and NIR studies (Sabbi et al., 2008; Weisz et al., 2008; Okamoto et al., 2015). The following sections present the UV diagnosis in both pictorial and graphic manners.
@@ III.2.2 Color-Color Diagram
From the color-color diagram, we can see that the FUV clusters and the diffuse regions inside the elliptical aperture possess the same regime following the model curve well;999There are two clusters with behaving as the exception. They are separately located on the galactic periphery and suspected to be background sources. the main body of the galaxy as a whole and the H-emission area are positioned identically at mag and mag, corresponding to the age Myr by reference to the model, as an average of the subregions. The overlaps between the different types of regions in this diagram imply all of them indistinguishable in stellar population. Therefore, the FUV clusters are not definitely younger than the diffuse regions but merely more in the number of stars, so is the H-emission area. In this situation, all of stars in Holmberg IX are supposed to be the same population born from a single SF event Myr ago, as also depicted by the single-peak Gaussian profiles in Figure 2. The recent occurrence of the global SF coincides with the hypothesis of the tidal origin for this galaxy. Combining with the above 2D map, we can demarcate for the galactic mainbody the lower age limit Myr that is newly discovered through the UV diagnosis in this work, and the upper limit Myr that is coincident with the previous optical-NIR studies ( Myr, Sabbi et al., 2008; Weisz et al., 2008; Okamoto et al., 2015). The lack of younger stellar populations suggests a very short duration for the past SF event. The model curve in Figure 3 is reproduced with the SFH of an instantaneous SF burst. Justification for the applied SFH is presented in Appendix A with more complex SFHs compared and discussed.
@@ III.3 Sparse Ionized Gas Revealed by H Mapping
For the purpose of exploring ionized gas linking current SF activities, we observed Holmberg IX through H channel in both of the direct narrowband imaging and FPI scanning modes (described in Section II.1). From a general view provided by the direct narrowband image, the H emission emerges dim, diffuse, and dispersive all over the galaxy. In the deep-exposure FPI map, it is spatially resolved into pieces of scattered spots and filaments located at northwest and southeast peripheries; most of the galactic body is left blank. The H appearance is opposite to the conspicuous and concentrative UV emission exhibited in the GALEX and Swift-UVOT images. Similar H-UV spatial contrast has been found in some other nearby dwarf galaxies and interpreted with SF propagation from galactic centers to outskirts, so that, after a certain timescale, current SF with strong H arrives in rims surrounding recently formed stellar populations bright in UV (Egorov et al., 2018; Gerasimov et al., 2024; Yarovova et al., 2024). Notwithstanding, in this work, such a behavior is observed in a TDG (candidate) for the first time.
@@
IV IMPLICATION
Under average conditions, weak H emission indicates a currently quiescent SF state. For Holmberg IX, this feature in combination with the UV diagnosis in Section III.2 potentially manifests global SF already ceased in most areas of the galaxy Myr ago. In this section, we assess current SF activities reflected by the marginal H plus the luxuriant FUV emissions, and deduce possible evolutionary pathways for this galaxy.
@@ IV.1 Current Star Formation State
The differences between the H and UV emissions in spatial position and radiative intensity inevitably cause considerable discrepancy between SFRs estimated from H and FUV integrated luminosities. By taking the aperture photometry and adopting the calibrations in K98, we obtain SFR(H) = () (corresponding to () ) and SFR(FUV) = () (corresponding to () ). Via specifically sampling the H-emission area, the selective photometry plus the K98 calibrations yields SFR(H) = () (corresponding to () ), SFR(FUV) = () (corresponding to () ), which narrows the gap between SFR(H) and SFR(FUV) effectively but remains insufficient. This examination demonstrates that the spatial mismatch between H and FUV emissions contributes a major part of the common discrepancy between SFR(H) and SFR(FUV) if galaxies are measured on an integrated basis. In this instance, different sources with H and FUV are actually sampled in practice. The selective photometry enables approaching the same H and FUV positions but still fails to totally fill the gap between the SFRs.
Different SF timescales traced by H and FUV luminosities are considered primarily responsible for SFR(H) lower than SFR(FUV), so-called ”H-to-FUV deficit”, no matter whether the two emissions are spatially displaced or identically located. From this perspective, SFR(H) takes an average of the latest Myr while SFR(FUV) averages past Myr (K98). In this case, SFR(FUV) higher than SFR(H) is a plausible consequence of different SF states tens of Myr ago (active) and at present (quiescent).
In order to quantify SF activity levels, we adopt the birthrate parameter defined as the ratio of current to mean past SFRs ( ; Kennicutt et al., 1994). The proportion valued by weighs the current star-formation level on the whole lifetime scale of a galaxy and reflects the evolution of SF activities. In this work, we set the mean past SFR equal to , where is total stellar mass, and is the stellar population age.101010The standard definition of the birthrate parameter contains a fraction of stellar mass returned to the galaxy in a stellar generation (Kennicutt et al., 1994). In this work, we assume a zero return fraction of stellar mass. Therefore, all mass is newly generated in every SF event, so that the birthrate parameter reaches the maximum. and are obtained with the stellar population synthesis model as best-fit parameters. The numerator SFRcurrent in is assigned as SFR(H) and SFR(FUV), respectively, which consequently brings on the birthrate parameter representing the SF state in recent Myr by and that in recent Myr by .
@@ IV.2 Subsequent Evolution and Ultimate Fate
The global SF dearth albeit the overall infancy for stellar populations requires rapid SF quenching, for which the reason seems to be a conundrum, particularly in view that the galaxy resides in a gas-rich environment ( and for Holmberg IX, presented in Sabbi et al., 2008). At the same time, the currently quiescent state conflicts with the commonly accepted requisite of continuous SF for TDGs (Duc & Mirabel, 1999, and references therein) and raises a question on subsequent evolution and ultimate fate of Holmberg IX. Without SF continuing, Holmberg IX is supposed to stay in a transient status at present and be transforming into a long-lived dwarf spheroidal galaxy due to gravitational contraction similar to those prevailing in the Local Group,111111In the Local Group, although there is no TDG, spheroidal galaxies are likely to originate from tidal events, which has been predicted by a number of numerical simulations (e.g., Mateo, 1998; Casas et al., 2012; Yang et al., 2014)., or to the opposite, dissolving into an ultra-diffuse galaxy (UDG) or intergalactic stellar stream because of centrifugal effects (and perhaps vanish in the Universe eventually). At this present moment, we are probably witnessing a rarely captured phenomenon for a TDG.
@@ V DISCUSSION
The above assessment and deduction rely on a high-mass SF scenario, which ascribes weak H but strong FUV emanations (or H-to-FUV deficit) to SFR low in recent Myr but high tens of Myr ago with bursty SF. In this case, aging of stellar populations leads to lack of high-mass, short-lifetime ionizing stars, which weakens H but sustains longer-time luminescence of FUV (Sullivan et al., 2000; Bell & Kennicutt, 2001; Iglesias-Páramo et al., 2004; Sullivan et al., 2004; Weisz et al., 2012). However, some observations of dwarf galaxies by resolving individual stars have suggested a more complicated situation that high-mass ionizing stars are not always absent for the H-to-FUV deficit (Weisz et al., 2008; McQuinn et al., 2009, 2010a, 2010b). In practical studies, IMF is usually assumed in a universal form with massive SF dominating, but if it turns out to be a variant in some circumstance, the speculation about galactic properties would be different. Besides IMF forms, a variety of other factors including SF modalities, interstellar medium conditions, observational bias, etc. are also able to interpret the H-to-FUV deficit recognized in a number of dwarf galaxies and extended outer disks (so-called extended UV disks, or XUV disks) of spiral galaxies (e.g., Hassani et al., 2024, and references cited in this section).
Instead of currently quiescent SF activities, the alternative explanations of H-to-FUV deficit allow the current occurrence of intense SF in agreement with the common expectation for TDGs. For unity of SFR(H) and SFR(FUV), Lee et al. (2009, hereafter denoted as L09) have recalibrated to SFR for dwarf galaxies with erg s-1. By applying the L09 calibration to Holmberg IX in our work, we obtain SFR(H) = () (corresponding to () ) for the aperture photometry, indeed reducing the offset from SFR(FUV), and SFR(H) = () (corresponding to () ) for the selective photometry, even exceeding SFR(FUV). Nevertheless, applicability of the L09 prescription to all dwarf galaxies is uncertain, and the resultant current SF activity for Holmberg IX is questionable until solid evidence is acquired. In below subsections, we elucidate all the other attributions of the H-to-FUV deficit and discuss the suitability for Holmberg IX.
@@ V.1 Low-mass or/and Stochastic SF
A low H amount is most directly ascribed to inherent shortage of high-mass ionizing stars. In addition to aging of stellar populations formed on a short or instantaneous timescale as suggested above, continuous SF with IMF in a low-mass dominant form is another mechanism innately preventing massive SF and consequently inducing H-to-FUV deficit. The variant IMF has been not only reproduced by stellar population synthesis modeling (Kroupa & Weidner, 2003; Weidner & Kroupa, 2005, 2006; Pflamm-Altenburg et al., 2007, 2009) but also observed in local galaxies with low surface brightness (Hoversten & Glazebrook, 2008; Meurer et al., 2009) and galactic XUV disks (Bruzzese et al., 2020; Rautio et al., 2024); still, it is not regarded universally effective (Hunter et al., 2010).
Even if stars keep forming with invariantly normal IMF, a lack of high-mass stars is also likely caused by stochasticity in SF events (da Silva et al., 2012, 2014). It is possible for random sampling of constant IMF at low SFR to undersample the number of high-mass ionizing stars (L09, and references therein), but the stochastic SF seems, still, not to suffice for the explanation either (Weisz et al., 2012). Therefore, two or more combinative effects of bursty, low-mass, or/and stochastic SF are considered more reasonable for restricting massive SF and more realistic for explaining observational H-to-FUV deficit (Fumagalli et al., 2011; da Silva et al., 2014; Lee et al., 2016). In this case, Holmberg IX is likely to keep forming stars for maintaining long-time survival, which hints there are possibly TDGs of the same type that failed to be recognized in the Universe.
@@ V.2 Loss of Ionizing Photons
If high-mass ionizing stars are always sufficient, H-to-FUV deficit is likely caused by loss of ionizing photons. The most probable mechanism is leakage of Lyman continuum in thin, porous, or/and (semi)transparent interstellar medium, which tends to decrease SFR(H) by up to 25% at fixed SFR(FUV) (Relaño et al., 2012). This mechanism has been applied to dwarf galaxies and XUV disks as an alternative explanation of H-to-FUV deficit (Hunter et al., 2010, 2011, 2013). However, it seems not possible for Holmberg IX located in a gas-rich environment where interstellar medium is rigorous and tight. Even though the HII regions are leaky and ionizing photons escape from natal areas, most of these photons still stay within the galaxy and supply diffuse emission lines, which is accounted for in integrated measurements.
@@ V.3 Underestimates of Existing H Emission
In addition to lack of ionizing stars and loss of ionizing photons, underestimates of H emission potentially contribute to the offset between observed and expected H levels, particularly for low surface-brightness systems (Hunter et al., 2010). However, much deeper imaging of dwarf galaxies than previous surveys has evinced that the unaccounted part of H emission (if any) compensates for only 5% flux at maximum and thus fails to suffice for a self-contained explanation (Lee et al., 2016). Nay, our ultra-deep FPI observations (with the sensitivity limit lower than that in Lee et al., 2016) and the net emission-line extraction from the 3D datacube guarantee the completeness of sampling H emission and thereof prevent the underestimation.
@@ VI SUMMARY
Via new observations deeply mapping H emission in Holmberg IX and combinatory analysis of archival/published data, we reveal a series of unique characteristics for this TDG candidate, distinct from other TDGs ever known. In the results, the galaxy appears extremely dustless, disproportionate to its rich gas and metal, with globally young stellar populations but currently inactive SF reflected by H-to-FUV deficit. In the galactic mainbody, all stars are suggested to be born from a bursty SF event Myr ago, which then rapidly ceased, demarcating the lower age limit Myr; younger stars reside in outskirts where mild SF remains as an aftermath of SF radial propagation. The recent quenching seems like a conundrum in view of the gas-rich location of the galaxy.
On a normal massive-SF assumption, the present quiescence conflicts with the common expectation for TDGs hosting continuous SF, which hints Holmberg IX currently staying in a rarely captured stage transforming into a dwarf spheroidal galaxy or oppositely into a(n) (ultra-)diffuse (intergalactic) system, which will probably dissolve in the end. Alternatively, by assuming a peculiar form of IMF, it is also possible for Holmberg IX to keep forming low-mass stars with H-to-FUV deficit and hence be able to maintain the current status for long-time survival. In whichever situation in reality, differing from conventional understandings, Holmberg IX deserves further investigations of its past experience, present state, and future destiny, as well as explorations of analogous TDGs in the Universe.